Planned downtime for firewall work, 29th October 2015

We will take our firewall offline intermittently from around 23:30 for essential maintenance following today's outages. Services should be normal after about 2 hours.

Service interruption to London data centre

29th October 2015 01:18 - 10:09
The service interruption to our London data centre this morning was caused by a power supply failure to a number of network elements. As soon as the power supply issue was resolved connection to the servers was re-established. We will soon be... read more
 

C1-Web-10 temporary monitoring failure

We received alerts that C1-Web-10 web server was offline, but the sites running on it were not reporting any failures. It appears the monitoring software itself had failed to re-start properly after a reboot for Windows updates. A re-start of the... read more
 

Reboot of Web-12 planned for early 22nd July 2015

We will be re-starting Web-12 to assign more CPU power early in the morning of 22nd July 2015.

Mapping services interruption early morning of 20th July

Some of our mapping services were compromised this morning, either responding very slowly or not at all. It appears that insufficient memory was allocated to the server following a restart for Windows updates. A manual restart of the server resolved ... read more
 

Advance Warning: Short service interruption to some website Wednesday 18th March 2015 at approx. 11pm.

Routine maintenance of two web servers means there will be a short interruption to the following websites at approx. 11pm on Wednesday 18th March 2015. The maintenance is expected to take approximately 15 minutes: Angling Diary website BASC... read more
 

Advanced Warning: Server reboots expected early morning Saturday 28th Feb. following Microsoft Updates

There are a number of critical Windows Updates that will applied to our servers early morning Saturday 28th Feb. (approx. 3-5am). This will cause some of the servers to reboot and will therefore cause a short period of service interruption. read more
 

Access to the SHINE website restored

Access to the SHINE website has been fully restored.

Services from ESDM-WEB-07 restored

All sites and services from ESDM-WEB-07 have been restored apart from SHINE (awaiting action by 3rd party). Affected services have been moved to another web server (C1-WEB-HBSMRGW) while we investigate the issues. The ESDM-WEB-07 server has... read more
 

Downtime affecting sites on ESDM-WEB-7

An unknown issue is preventing ESDM-WEB-7 from connecting to the network after a restart. The following sites/systems are affected: Documentation sites: Hbsmrdocumentation.esdm.co.uk Librarylinkdocs.esdm.co.uk Maplinkdocs.esdm.co.uk... read more
 
Page 7 of 7 << < 1 2 3 4 5 6 7
Find out more